OpenStudy (anonymous):

what is 6xsquared-24x+24/x-4

OpenStudy (anonymous):

and what is the domain of this

OpenStudy (anonymous):

\[6x^{2} - \frac{24}{x} - 4\]This? is it equal to 0?

OpenStudy (anonymous):

yea its that question and n0

OpenStudy (anonymous):

Well if it's not equal to 0, there's nothing to solve for. As for the domain, the only restriction it has is from the x in the denominator of 24/x. This means that x can not be 0 because it would be dividing by 0. So the domain is everything but 0: \[D: (-\infty, 0)\cup(0, \infty)\]